The first test in Introductory Algebra had a mean of 82 with a standard deviation of 15. If a student made a 94 on the test, what is the corresponding z-score?
a) -0.58 b) -0.47 c) 0.47 d) 0.80 e) NOT

Answer by fractalier(6550)

The z score is found by
z = (x - mean)/sigma
z = (94-82)/15
z = 12/15 = .8
Choice d.